# Toward a Vision-Language Foundation Model for Medical Data: Multimodal Dataset and Benchmarks for Vietnamese PET/CT Report Generation

### v1

@@ Abstract
Vision-Language Foundation Models (VLMs), trained on large-scale multimodal datasets, have driven significant advances in Artificial Intelligence by enabling rich cross-modal reasoning. Despite their success in general domains, applying these models to medical imaging remains challenging due to the limited availability of diverse imaging modalities and multilingual clinical data. Most existing medical VLMs are trained on a subset of imaging modalities and focus primarily on high-resource languages, thus limiting their generalizability and clinical utility. To address these limitations, we introduce a novel Vietnamese-language multimodal medical dataset comprising 1,567,062 paired CT-PET images and corresponding 2,757 full-length clinical reports. This dataset is designed to fill two pressing gaps in medical AI development: (1) the lack of PET/CT imaging data in existing VLMs training corpora, which hinders the development of models capable of handling functional imaging tasks; and (2) the underrepresentation of low-resource languages, particularly the Vietnamese language, in medical vision-language research. To the best of our knowledge, this is the first dataset to provide comprehensive PET/CT-report pairs in Vietnamese. We further introduce a training framework to enhance VLMs’ learning, including data augmentation and expert-validated test sets. We conduct comprehensive experiments benchmarking state-of-the-art VLMs on downstream tasks, including medical report generation and visual question answering. The experimental results show that incorporating our dataset significantly improves the performance of existing VLMs. However, despite these advancements, the models still underperform on clinically critical criteria, particularly the diagnosis of lung cancer, indicating substantial room for future improvement. We believe this dataset and benchmark will serve as a pivotal step in advancing the development of more robust VLMs for medical imaging, particularly in low-resource languages, and improving their clinical relevance in Vietnamese healthcare.
@@ 1 Introduction
Vision-Language Foundation Models (VLMs) have rapidly evolved as a cornerstone of modern Artificial Intelligence (AI), capable of jointly modeling information across visual and textual modalities. These models are typically pre-trained on diverse datasets encompassing billions of image-text pairs, enabling them to acquire generalized and transferable representations [1, 2, 3, 4, 5]. This cross-modal alignment allows VLMs to bridge the semantic gap between images and language, facilitating downstream tasks such as image captioning [6, 7], visual question answering [4, 5, 6, 7], report generation [8, 9], and even zero-shot image classification [1, 10], with minimal task-specific supervision.
@@
    We introduce a comprehensive dataset comprising 1,567,062 paired PET/CT images along with 2,757 full-length Vietnamese clinical reports. The dataset spans a demographically and pathologically diverse patient population, reflecting real-world clinical variability. It provides a valuable resource for advancing the training of medical VLMs, with the potential to support a broader range of modalities and enable multilingual development, especially for low-resource languages.
@@
Dataset Name
Text Image Modality Size
@@ 2.1 Existing Medical Multimodal Datasets
To overcome current limitations, we present ViMed-PET, a comprehensive PET/CT image-report dataset composed of two main parts, as summarized in Table 3. The first part consists of the original dataset, including paired PET/CT images and their corresponding clinical reports, collected directly from hospitals (Table 3a). The second part comprises a series of augmented datasets derived from the original data. These augmentations aim to increase the diversity and richness of the dataset while enabling more fine-grained alignment between the visual and textual modalities (Table 3b). The following sections detail the structure of ViMed-PET and the methodology used to construct it.
@@ 2.2 Data Description
The proposed ViMed-PET dataset is sourced exclusively from a reputable hospital in Vietnam, ensuring authenticity and reliability. It includes whole-body PET/CT images paired with corresponding medical reports, comprising 2,757 studies collected over four years. Note that the dataset does not contain complete data for all 12 months of each year, as detailed in the statistics shown in Table 3. In total, the dataset contains 1,567,062 paired CT-PET images and 2,757 corresponding full-length clinical reports. Each study includes approximately 250–500 paired CT and PET slices, covering the area from the head to the upper thighs (just above the knees). The dataset encompasses various disease cases such as lung cancer, thyroid cancer, and other conditions, representing a broad range of clinical scenarios. The images are stored in the Digital Imaging and Communications in Medicine (DICOM) format, including pixel data and relevant metadata such as patient age, sex, body weight, radiotracer activity, and other acquisition parameters. Acquired using GE Discovery 710 PET/CT and GE Discovery STE PET/CT systems, the images provide high-quality data for analysis. Furthermore, the PET images have undergone attenuation correction using the corresponding CT data to ensure accurate representation. Each medical report, stored in DOCX format, corresponds to a single PET/CT study and contains detailed patient information, clinical status, medical history, scanning methods, and physician observations. This makes the dataset rich in both imaging and clinical information.
@@ 2.3 Data Pre-processing
Body part-based data partition. To expand the dataset and improve the data quality used for training and evaluating the model, we divide each study, comprising a PET/CT image and its corresponding report, into three anatomically distinct regions: head-neck, chest, and abdomen-pelvis. This approach results in a total of 8,271 paired image-report samples. To preserve continuity across adjacent regions and avoid the loss of critical contextual information at segment boundaries, we introduce a 20-slice overlap between neighboring segments. Specifically, the head-neck region corresponds to the first 20% of the body scan, the chest segment begins 15% below the last slice of the head-neck region, and the abdomen-pelvis segment includes the remainder of the scan from the chest to the pelvis. This segmentation strategy increases the number of training samples and significantly improves the alignment between the visual and textual modalities. By localizing image content and clinical descriptions to specific anatomical regions, we enable more precise fine-tuning of models, ultimately enhancing their ability to learn region-specific patterns and improving overall performance.
@@ 2.4 Dataset Construction
Visual question answering dataset. The VQA dataset aims to fine-tune VLMs by enabling context-aware, multi-turn dialogue about biomedical images for tasks like diagnostic reasoning and clinical decision-making. The dataset is composed of two parts: single-turn and multi-turn conversations. First, we created a set of 27,855 image-related questions that prompt descriptive answers. These questions were then randomly sampled and paired with each PET/CT image-report pair from the original dataset to generate single-turn VQA samples. Next, we employ the few-shot prompting strategy to guide the GPT-4o [12] in generating multi-turn conversations based on the clinical reports from the original dataset (detailed in the Appendix B.1). In total, we constructed 8,271 multi-turn conversations as part of this dataset.
@@ 4.2 PET/CT VQA Task
Cosmos
Tokenizer
Mistral-7B 28.09 62.92 48.37 79.25
@@ 5 Conclusion
In this study, we introduced ViMed-PET, a high-quality dataset comprising 1,567,062 paired PET/CT images and 2,757 Vietnamese clinical reports, covering a wide range of patient cases. We also developed a clinically validated lung cancer test set to support meaningful evaluation beyond conventional NLP metrics. Additionally, we proposed a data augmentation strategy that enhances both visual and textual inputs to improve the fine-tuning of medical vision-language models. Models fine-tuned on ViMed-PET demonstrated substantial gains, achieving improvements in both standard NLP metrics and clinical evaluation scores compared to pretrained baselines.
Limitations and Societal Impacts. We acknowledge that clinical reports often follow standardized formats, which can limit output diversity. Clinical results with F1 scores around 50% highlight the challenges in modeling PET/CT data and performing complex medical reasoning. Although the proposed ViMed-PET dataset contains PET/CT and CT images, our benchmark focuses exclusively on PET/CT image. This decision is based on the observation that report content is primarily driven by PET information, with minimal reference to CT anatomical details. In future work, we plan to extend our approach to better incorporate CT data and further improve the accuracy of VLMs. This study provides a foundation for such enhancements. We believe ViMed-PET serves as a valuable resource for advancing medical vision-language modeling in the low-resource Vietnamese language and the underexplored PET/CT domain, supporting more equitable AI development in healthcare.
Acknowledgments and Disclosure of Funding
This research is funded by Hanoi University of Science and Technology (HUST) under grant number T2024-TĐ-002. We would like to thank the NVIDIA Academic Grant Program 2025 for providing the computing resources used in this research.
Appendix A Medical Vision-Language Models
Appendix B Technical Appendices
B.1 Visual Question Answering Dataset
B.2 Study Comparison Dataset
B.3 Training and Model Configurations
B.3.1 Fine-tuning Vision Encoders
B.3.2 Fine-tuning VLMs
B.3.3 Fine-tuning Resources
      Model
Computational Resources
Concept Alignment LoRA Fine-tuning
Vision Language Time (Hours) Mem (GB) Time (Hours) Mem (GB)
Fine-tuned
CT-ViT Mistral-7B 2.00 61.0 12.00 76.0
LLaMA-2-7B 2.00 62.0 12.00 73.0
Cosmos Tokenizer Mistral-7B 1.83 47.5 11.00 70.0
LLaMA-2-7B 1.75 46.5 11.00 71.6
B.4 Clinical Evaluation Metrics
  • FDG: {increase, not increase}
  • Position: {mediastinum, lung, abdomen, axilla, cervical region}
B.5 Task Evaluation by GPT-4o
Appendix C Additional Results
C.1 PET/CT VQA Task
Model Settings NLP Metrics $\uparrow$
Vision Language O G C BLEU-4 R-1 R-L BERT
Baseline
LLaVA-Med [27]   3.39 47.83 33.82 75.86
M3D [28]   0.03 11.80   9.66 59.87
RadFM [29]   0.04 11.71 12.24 61.93
GPT-4o* [12]   3.01 49.35 30.09 71.92
Fine-tuned
CT-ViT
23.22 57.61 43.80 77.06
Mistral-7B 31.33 65.61 51.22 82.05
31.14 65.10 50.33 81.80
26.93 56.01 42.28 75.31
LLaMA-2-7B 26.36 56.48 42.79 77.73
31.36 59.14 48.00 76.72
Cosmos
Tokenizer
20.01 58.17 42.54 76.47
Mistral-7B 25.71 61.05 46.59 78.49
28.09 62.92 48.37 79.25
25.83 61.80 46.58 78.87
LLaMA-2-7B 26.11 62.26 47.05 79.39
28.40 63.29 48.76 79.35
C.2 Report Generation and VQA Samples

### v3

@@ Abstract
Vision-Language Foundation Models (VLMs), trained on large-scale multimodal datasets, have driven significant advances in Artificial Intelligence (AI) by enabling rich cross-modal reasoning. Despite their success in general domains, applying these models to medical imaging remains challenging due to the limited availability of diverse imaging modalities and multilingual clinical data. Most existing medical VLMs are trained on a subset of imaging modalities and focus primarily on high-resource languages, thus limiting their generalizability and clinical utility. To address these limitations, we introduce a novel Vietnamese-language multimodal medical dataset consisting of 2,7572{,}757 whole-body PET/CT volumes from independent patients and their corresponding full-length clinical reports. This dataset is designed to fill two pressing gaps in medical AI development: (1) the lack of PET/CT imaging data in existing VLMs training corpora, which hinders the development of models capable of handling functional imaging tasks; and (2) the underrepresentation of low-resource languages, particularly the Vietnamese language, in medical vision-language research. To the best of our knowledge, this is the first dataset to provide comprehensive PET/CT-report pairs in Vietnamese. We further introduce a training framework to enhance VLMs’ learning, including data augmentation and expert-validated test sets. We conduct comprehensive experiments benchmarking state-of-the-art VLMs on downstream tasks, including medical report generation and visual question answering. The experimental results show that incorporating our dataset significantly improves the performance of existing VLMs. However, despite these advancements, the models still underperform on clinically critical criteria, particularly the diagnosis of lung cancer, indicating substantial room for future improvement. We believe this dataset and benchmark will serve as a pivotal step in advancing the development of more robust VLMs for medical imaging, particularly in low-resource languages, and improving their clinical relevance in Vietnamese healthcare.
@@ 1 Introduction
Vision-Language Foundation Models (VLMs) have rapidly evolved as a cornerstone of modern Artificial Intelligence (AI), capable of jointly modeling information across visual and textual modalities. These models are typically pre-trained on diverse datasets encompassing billions of image-text pairs, enabling them to acquire generalized and transferable representations [41, 56, 23, 7]. This cross-modal alignment allows VLMs to bridge the semantic gap between images and language, facilitating downstream tasks such as image captioning [6, 28], visual question answering [23, 7, 6, 28], report generation [54, 40], and even zero-shot image classification [41, 18], with minimal task-specific supervision.
@@
    We introduce a comprehensive dataset comprising 2,757 whole-body PET/CT volumes from independent patients along with full-length Vietnamese clinical reports. The dataset spans a demographically and pathologically diverse patient population, reflecting real-world clinical variability. It provides a valuable resource for advancing the training of medical VLMs, with the potential to support a broader range of modalities and enable multilingual development, especially for low-resource languages.
@@
Dataset Name Text Image Modality Size
@@
For our ViMed-PET dataset, 1,567{1,567}K paired slices correspond to 2,757 paired whole-body PET/CT volumes.
@@ 2.1 Existing Medical Multimodal Datasets
To overcome current limitations, we present ViMed-PET, a comprehensive PET/CT image-report dataset composed of two main parts, as summarized in Table 3. The first part consists of the original dataset, including paired PET/CT images and their corresponding clinical reports, collected directly from a hospital (Table 3a). The second part comprises a series of augmented datasets derived from the original data. These augmentations aim to increase the diversity and richness of the dataset while enabling more fine-grained alignment between the visual and textual modalities (Table 3b). The following sections detail the structure of ViMed-PET and the methodology used to construct it.
@@ 2.2 Data Description
The proposed ViMed-PET dataset is collected exclusively from a national tertiary general hospital in Vietnam, one of the country’s largest medical centers. As a high-volume referral institution receiving patients from across all regions, its data reflect broad clinical diversity and ensure high representativeness and reliability. It consists of 2,757 paired CT-PET volumes (equivalent to 1,567,062 paired CT-PET slices) collected over four years, each accompanied by a corresponding full-length clinical report. Note that the dataset does not contain complete data for all 12 months of each year, as detailed in the statistics shown in Table 3. Each study includes approximately 250–500 paired CT and PET slices, covering the area from the head to the upper thighs (just above the knees). The dataset encompasses various disease cases such as lung cancer, thyroid cancer, and other conditions, representing a broad range of clinical scenarios. The images are stored in the Digital Imaging and Communications in Medicine (DICOM) format, including pixel data and relevant metadata such as patient age, sex, body weight, radiotracer activity, and other acquisition parameters. Acquired using GE Discovery 710 PET/CT and GE Discovery STE PET/CT systems, the images provide high-quality data for analysis. Furthermore, the PET images have undergone attenuation correction using the corresponding CT data to ensure accurate representation. Each medical report, stored in DOCX format, corresponds to a single PET/CT study and contains detailed patient information, clinical status, medical history, scanning methods, and physician observations, making the dataset rich in both imaging and clinical information. All data are obtained under the oversight of an Institutional Ethics Committee (Ethics Approval No. 6184/CN-HDDDBV). Informed consent for anonymized research use is obtained by the hospital, in accordance with institutional policy and national regulations.
@@ 2.3 Data Pre-processing
Body part-based data partition. To expand the dataset and improve the data quality used for training and evaluating the model, we divide each study, comprising a PET/CT volume and its corresponding report, into three anatomically distinct regions: head-neck, chest, and abdomen-pelvis. This approach results in a total of 8,271 paired image-report samples. To preserve continuity across adjacent regions and avoid the loss of critical contextual information at segment boundaries, we introduce a 20-slice overlap between neighboring segments. The proportional boundaries of each region are adaptively determined based on the patient’s body height, ensuring flexibility across subjects of different statures. The specific ratios for region division follow expert guidelines, with the head–neck typically covering approximately the first 20% of the body length, the chest starting around 15% below the last slice of the head–neck region, and the abdomen–pelvis including the remainder from the chest down to the pelvis. This segmentation strategy increases the number of training samples and significantly improves the alignment between the visual and textual modalities. By localizing image content and clinical descriptions to specific anatomical regions, we enable more precise fine-tuning of models, ultimately enhancing their ability to learn region-specific patterns and improving overall performance.
@@ 2.4 Dataset Construction
Visual question answering dataset. The VQA dataset aims to fine-tune VLMs by enabling context-aware, multi-turn dialogue about biomedical images for tasks like diagnostic reasoning and clinical decision-making. The dataset is composed of two parts: single-turn and multi-turn conversations. First, we created a set of 27,855 image-related questions that prompt descriptive answers. These questions were then randomly sampled and paired with each PET/CT image-report pair from the original dataset to generate single-turn VQA samples. Next, we employ the few-shot prompting strategy to guide the GPT-4o [17] in generating multi-turn conversations based on the clinical reports from the original dataset (detailed in the Appendix B.1). In total, we construct 8,271 multi-turn conversations as part of this dataset.
@@ 4.2 PET/CT VQA Task
Cosmos Tokenizer Mistral-7B 28.09 62.92 48.37 79.25
@@ 5 Conclusion
In this study, we introduced ViMed-PET, a high-quality dataset comprising 2,757 paired whole-body PET/CT volumes and 2,757 Vietnamese clinical reports, covering a wide range of patient cases. We also developed a clinically validated lung cancer test set to support meaningful evaluation beyond conventional NLP metrics. Additionally, we proposed a data augmentation strategy that enhances both visual and textual inputs to improve the fine-tuning of medical vision-language models. Models fine-tuned on ViMed-PET demonstrated substantial gains, achieving improvements in both standard NLP metrics and clinical evaluation scores compared to pretrained baselines.
Limitations and Societal Impacts. We acknowledge that clinical reports often follow standardized formats, which can limit output diversity. Clinical results with F1 scores around 50% highlight the challenges in modeling PET/CT data and performing complex medical reasoning. Although the proposed ViMed-PET dataset contains PET and CT volumes, our benchmark focuses exclusively on PET/CT imaging. This decision is based on the observation that report content is primarily driven by PET information, with minimal reference to CT anatomical details. In future work, we plan to extend our approach to better incorporate CT data and further improve the accuracy of VLMs. This study provides a foundation for such enhancements. We believe ViMed-PET serves as a valuable resource for advancing medical vision-language modeling in the low-resource Vietnamese language and the underexplored PET/CT domain, supporting more equitable AI development in healthcare.
@@ 80. Guidelines:
Appendix A Medical Vision-Language Models
Appendix B Technical Appendices
B.1 Visual Question Answering Dataset
B.2 Study Comparison Dataset
B.3 Training and Model Configurations
B.3.1 Fine-tuning Vision Encoders
The original Cosmos Tokenizer uses causal attention for video tasks, modeling forward-only temporal relationships. However, 3D PET volumes have bidirectional spatial relationships, making causal attention less appropriate. To assess this, we compared both settings (with vs. without causal masking) using LLaMA-2-7B on the ViMed-PET dataset across two tasks: Report Generation and Visual Question Answering (VQA). As shown in Table 6, removing causal attention yields significantly better performance across both tasks.
Table 6: Performance comparison of causal attention settings on ViMed-PET across two tasks: Report Generation and Visual Question Answering (VQA). R-1 and R-L denote ROUGE-1 and ROUGE-L scores. $\uparrow$ indicates higher is better.
Model Causal Attention Report Generation VQA
[…]
Cosmos Tokenizer LLaMA-2-7B 54.99 77.82 68.88 87.86 19.87 55.60 41.10 76.49
$\times$ 57.59 79.26 73.91 90.35 28.40 63.29 48.76 79.35
B.3.2 Fine-tuning VLMs
B.3.3 Fine-tuning Resources
      Model Computational Resources
Concept Alignment LoRA Fine-tuning
Vision Language Time (Hours) Mem (GB) Time (Hours) Mem (GB)
Fine-tuned CT-ViT Mistral-7B 2.00 61.0 12.00 76.0
LLaMA-2-7B 2.00 62.0 12.00 73.0
Cosmos Tokenizer Mistral-7B 1.83 47.5 11.00 70.0
LLaMA-2-7B 1.75 46.5 11.00 71.6
B.4 Clinical Evaluation Metrics
  • FDG: {increase, not increase}
  • Position: {mediastinum, lung, abdomen, axilla, cervical region}
B.5 Task Evaluation by GPT-4o
Appendix C Additional Results
C.1 PET/CT VQA Task
Model Settings NLP Metrics $\uparrow$
Vision Language O G C BLEU-4 R-1 R-L BERT
Baseline LLaVA-Med [24]   3.39 47.83 33.82 75.86
M3D [5]   0.03 11.80   9.66 59.87
RadFM [50]   0.04 11.71 12.24 61.93
GPT-4o* [17]   3.01 49.35 30.09 71.92
Fine-tuned CT-ViT 23.22 57.61 43.80 77.06
Mistral-7B 31.33 65.61 51.22 82.05
31.14 65.10 50.33 81.80
26.93 56.01 42.28 75.31
LLaMA-2-7B 26.36 56.48 42.79 77.73
31.36 59.14 48.00 76.72
Cosmos Tokenizer 20.01 58.17 42.54 76.47
Mistral-7B 25.71 61.05 46.59 78.49
28.09 62.92 48.37 79.25
25.83 61.80 46.58 78.87
LLaMA-2-7B 26.11 62.26 47.05 79.39
28.40 63.29 48.76 79.35
C.2 Report Generation and VQA Samples